1. Richard e-mailed me that he and Monica differed about the impact of his extra spending...

1.

Richard e-mailed me that he and Monica differed about the impact of his extra spending over the past 15 years. He calculated it at about $3,000 a year. He said the total cost of $45,000 was well within his capability to make up. Monica said the cost was much higher and asked that they compute it. They had been offered an investment of $20,000 that would pay $70,000 in 20 years. They want to know if they should take it. Finally, Richard could sign up for an annuity at work. It would cost $100,000 at age 65 and provide payments of $8,000 per year over his expected 17-year life span. He wants to know if it is attractive. The appropriate market rate of return on investments is 7 percent after tax.

Calculate what the $3,000-per-year deficit, had it been invested, would have amounted to at the end of the 15-year period.

2. Calculate the return on the proposed $20,000 investment and indicate the factors entering into your recommendation to accept or reject it.

Choose TRUE if you accept it

Choose FALSE if you reject it.
Answer #1

1:

Value of $3000 after 15 years = A*((1+r)^n-1) /r

=3000*(1.07^15-1)/0.07

=75387.07

2: Value of $20,000 in 20 years at 7% rate

FV= PV*(1+r)^n

= 20000*1.07^20

=77393.69

However they are getting only 70,000 hence it should not be taken.

Choose False

3: FV of annuity = A*((1+r)^n-1) /r

=8000*(1.07^17-1)/0.07

=246721.7

However he is getting only 100,000. Hence it should not be taken.